# Gaia Data Release 3: spectroscopic binary-star orbital solutions

### v1

@@ 3 The underlying orbital models
Four kinds of orbital solutions are considered here: they are labelled SB1, SB1C, TrendSB1 and StochasticSB1. They are described here below. The single-line RVs treated here and consequently the deduced motion is intended to refer to the sole object that is visible in the spectra. The model will be fitted through a classical least-square procedure and the selected objective function is the -function except when otherwise stated. The uncertainties on the parameters are calculated in the classical way on the basis of the diagonal elements of the inverse of the curvature matrix (i.e. the variance-covariance matrix, see Bevington & Robinson 2003). Despite the fact that we use the Greek symbol for all (i.e. concerning input data as well as output parameters) uncertainties, it should be pointed out that all of them are of course estimators. They are always expressed as 1 values.
@@ 3.1 Orbital models of type SB1
where is the time of passage at periastron. The is the argument of periastron measured in the sense of the orbit with its origin at the ascending node. In this model, the eccentricity enters as a strongly non-linear parameter. This is also the case for . The parameters to be determined are , , , , , (respectively, the period, the centre-of-mass velocity, the semi-amplitude, the eccentricity, the argument of periastron and the time of passage at periastron). The first step of the computation is devoted to the determination of the period. Some significance tests are made on the very existence of the period. If the period is not significant, the model is fitting noise and is considered as being invalid. This is a statistical decision. Owing to the insufficient span of time covered for the DR3, no apsidal motion is considered; is considered constant in time.
@@ 3.3 Models of type TrendSB1
It can happen that the RV evolution as a function of time just exhibits a trend. This could be a transient behaviour (e.g. outburst) just observed by chance or correspond to a small piece of slowly accelerated motion (e.g. of a stellar atmosphere). It could also correspond to an orbital motion with a period well longer than the span of observing time covered by the satellite. This span of time is 34 months for the DR3. However, it could be slightly shorter for some objects because it combines with the gaps in the scanning law. The present modelisation is restricted to polynomials of degrees one to four and presenting only one extremum on the interval. The degree of the polynomial is chosen on the basis of an F-test for nested models. In practice, it turned out that only the one- and two-degree polynomials were necessary. No higher degree fit was required. The formalism used for the fit is the same as for Hipparcos (ESA 1997). The TrendSB1 solution is compared to the orbital solutions using a Fisher Snedecor F-test on the comparison of the respective s. It should be made clear that this comparison could be an approximation.
@@ 3.5 The downstream combination of solutions
The outputs of the NSS spectroscopic pipeline are delivered to the combiner that operates downstream. The task of the combiner is to inspect the output of the astrometric, photometric and spectroscopic NSS channels in order to detect possible solutions corresponding to the same object and that could be combined. This will be described in Sect. 7.2.
@@ 4 Determination of the periodicity
The first step of the analysis is the search for a periodic behaviour of the RVs. The analysis is based on a periodogram (actually a frequencygram), which has the nature of the Fourier Power Spectrum. Since the time sampling of the Gaia satellite is particularly complex, the classical Fourier formalism is not applicable. To well understand the limitation of the method, it is necessary to have some knowledge of the impact of the sampling and of a few characteristic timescales.
@@ 4.1 Effect of the Gaia sampling
In order to explain the effect of the Gaia sampling, we have to start from a single, idealistic, case and progressively add some complexity. We start by considering the spinning of the satellite on itself that makes that a star is measured every 0.25 d (this is our first basic idealistic case since this simple view is never occurring in practice). The existence of the sample periodicity concerns both fields of view individually. Therefore, 0.25 d is the typical, smallest regular and periodic timescale in the sampling. The 0.25 d step should generate in the time-series some aliasing with a typical 2 d-1 Nyquist frequency. This means that a progenitor frequency will be duplicated at , , , and so on (in d-1 units). The duplication is also extended to the negative frequencies. This duplication represents a pure aliasing, since the sampling is in this hypothetical case regular. Therefore, the aliasing is pure and the Nyquist frequency exists. In such a case, intermediate and large periods could have aliases towards periods of 0.25 to 0.5 d (and vice versa).
@@
However, the existence of two fields of view separated by 1065 on the sky renders the situation more complex by adding a regularity that is not properly speaking periodic. Thus, the aliasing is no more pure, we will call it pseudo-aliasing. The semi-amplitude spectral window is a tool that facilitates the analysis of such phenomenon. It has been introduced by Deeming (1975, 1976). In the case of an equidistant sampling, the spectral window is the well-known Dirac comb (or Shah function, see Bloomfield 1976). The general spectral window represents an extension of this Dirac comb to the general case of odd sampling. The property that the generalised Fourier Spectrum of the sampled data is the result of the convolution of the underlying Fourier Spectrum of the signal by the spectral window remains true for almost any kind of sampling (Deeming 1975, 1976). If the satellite was limited to a rotation on itself, the semi-amplitude spectral window corresponding to the combined two fields of view would have the aspect given in the upper panel of Fig. 1. Contrary to the classical case of pure aliasing, the aliasing is here not pure since all the peaks have different heights. Besides the main peak at , the next first peak is at d-1, but it is mildly high. The peak having a height comparable to the peak is the one at d-1 which could represent some almost pure aliasing. This figure represents an idealistic case. First, it is assumed that the data points in each field of view are perfectly regularly distributed. This is not the case because several phenomena induce a small perturbation of the time of the observations. The variations of the Basic Angle as well as micro-clanks and meteoritic-hits are perturbing the regularity (Gaia Collaboration et al. 2016b). We do not further consider these effects in our explanation. The satellite is not only spinning on itself, but also follows a more complicated motion. First, the rotation axis is precessing along an axis directed towards a fictitious nominal position of the Sun (compensating for the orbit of the satellite around the Sun-Earth Lagrangian point L2). In addition, the satellite located around the Lagrangian point L2 is orbiting around the Sun with the Earth annual motion (change in ecliptic longitude of the Sun, Gaia Collaboration et al. 2016b). The major consequence is that, for a given object, the majority of the turns of the spinning satellite will not correspond to a transit of the object and thus no observation will be acquired for this object. Most of the observations are not done and medium and large gaps occur during which the field of view is not directed towards the star. This effect is dependent on the position of the object on the sky (Gaia Collaboration et al. 2016b). Therefore, the sky sampling is markedly non-uniform (see section 5.2 of Gaia Collaboration et al. 2016b).
@@
Practically, this generates in the spectral window a power leakage from the main peak at d-1 towards other frequencies. This effect is illustrated in the other panels of Fig. 1. We produce typical time-series of Gaia RVS with the GOST tool222https://gaia.esac.esa.int/gost/. Fig. 1 second panel represents a typical spectral window of a time-series containing 26 observations, the third panel 19 and the bottom one 14 observations (against the observations for the first panel). This small number of observations induces an aliasing but the aliasing typical of the Gaia sampling is not leaking power towards a very few other frequencies but towards numerous frequencies. Besides the main peak, there is a leakage towards a forest of frequencies. The contrast between the parasitic peaks and the main one is decreasing with the number of data points (following a for the semi-amplitude). Typically, pseudo-aliasing can occur at any . If existing, the value of is most probably located between and let say 3 d-1. The structure due to the motion considered in the top panel of Fig. 1 almost totally disappeared. In the power spectrum of the signal, the combination of noise and of aliasing is going to generate parasitic peaks at frequencies that mainly depend on the position of the object on the sky. Consequently, the time sampling is complex, and the aliasing is dependent on the star being observed.
@@
Looking at the third panel of Fig. 1, one can notice marked aliasing at and at . Therefore, in a few cases, noise could conspire to shift the largest peak position from the progenitor peak by this amount, thus producing pseudo-aliases of the progenitor that could dominate. However, in the present case, predominant aliasing seems to be restricted to small , i.e. to small shifts of the progenitor peak. Other values of will be associated (or not) to other regions of the sky. Large shift aliasing is clearly possible but rather for values . This should have no impact in the framework of spectroscopic binaries.
@@
Here above, we have shown three examples illustrating the strong leakage characterising the Gaia sampling (for RVS). An alternative view would be to compute some kind of typical spectral window. We obtain such a result using simulations. We generated 65536 positions randomly and uniformly distributed over all the sky. For each position, we extracted from the GOST software the series of observational epochs typical of this particular position and corresponding to the Nominal Scanning Law (NSL) for RVS. On purpose, we discarded the first month of scanning which was characterised by the Ecliptic Pole Scanning Law (EPSL). For the latter, the precession of the satellite was stopped and the scanning law is more relevant to the scanning due to the spinning of the satellite except for a slow rotation due to the orbital motion around the Sun. We thus neglect the EPSL to concentrate on the effect of the NSL. For each simulation, we derive the corresponding amplitude spectral window (for the RVS instrument). Figure 2 illustrates (in red) the median of the amplitude spectral windows. Despite the fact that we restricted our simulations to the NSL, the aliasing due to the spinning of the satellite is clearly visible. However, even the peaks at d-1 (and at d-1) are rather low compared to the peak at d-1, at most 30% (in amplitude). We also plot in Fig. 2 (in black) 32 randomly selected amplitude spectral windows. Clearly, the dispersion of the various realisations of the amplitude spectral window is rather large. Also, near the possible alias frequencies, the amplitude spectral window could present values nearing zero. This confirms the fact that in most cases, the spinning pseudo-aliasing will have little, if any, impact on the global aliasing. Although a few positions on the sky could still exhibit the effect of the spinning aliasing. Of course, if, over the 34 months, the EPSL scanning is dominating the definition of the times in the analysed time-series, the effect will increase exhibiting a stronger spinning aliasing. We must insist here on the fact that no signal is analysed in the present section; the dispersion of the amplitude spectral windows is originating from the randomness of the selection of the sky positions and from the variety of nature of the aliasing as a function of the latter. For any time-series, the observing time could be used to pre-calculate the particular spectral window independently of the signal.
@@ 4.2 Search for the periodicity
When a periodicity is present in a time-series, the periodogram exhibits a peak at the corresponding frequency (the progenitor), but other peaks are also present at the location of the pseudo-aliases of the progenitor. Contrary to the case of even sampling, the actual sampling generates pseudo-aliases that are in principle lower than the progenitor peak. However, the combination of the aliasing with background noise or other phenomena (blending) induces the fact that the choice of the true progenitor among the aliased peaks is never fully secure.
@@
The total span of time of the DR3 time-series is at most of 34 months and thus the largest periodicity accessible from the point of view of the Fourier corresponds roughly to a frequency of 1/1000 d-1. This value could actually be larger taking into account the gaps generated by the actual sampling law. The constraints should thus be expressed as a function of the actual length of the time-series and thus as a function of . It has been decided that it would be adequate to limit the range of investigation for periodicities to since beyond the possible orbit is certainly not well determined. However, from the practical point of view, we decided to limit the range for period search to although no orbit are computed for periods located in this extension (see Sect. 6.3 for more details). The adopted frequency step is actually where is the actual total time-span of the individual time-series. This represents an oversampling of a factor one hundred that gives directly access to the amplitude of the signal as being the height of the peak. From the previous section discussion, we decided that the adopted largest frequency to be explored is 4 d-1. This choice is rather atypical since it is chosen to correspond to the spinning period.
@@
If the star indeed varies in a periodic way, the largest peak in the periodogram is an estimate of the possible period but definitively not a proof of its existence. One should consider that a periodicity is proven if a minimum of three identical cycles are observed. Although the long periods are indicative, they are associated to large errors due to the small number of cycles included in the computation.
@@
The method used is the one of Heck-Manfroid-Mersch (hereafter HMM: Heck et al. 1985; Gosset et al. 2001), which advantageously replaces the classical Fourier method (see e.g. Bloomfield 1976) in the case of odd sampling. In principle, the highest peak is assumed to be the progenitor one and to correspond to the orbital period. This last point is particularly true for the case of RV orbital motion, since most of the power is gathered in the fundamental frequency; the power of the peaks corresponding to high harmonics is gradually decreasing in the case of the Keplerian motion. This is due to the fact that RV curves always present one simple cycle during the orbital period. On the contrary, this is usually not the case for light curves of eclipsing binaries and ellipsoidal variables where most of the power is gathered in twice the orbital frequency, corresponding to the average depth of the minima, whereas the power at the orbital frequency is dominated by the difference between the primary and secondary minima and thus is much weaker. In their paper, HMM introduced the statistic which is a power related statistic (see Eq. A.11 in Gosset et al. 2001). At a predetermined frequency and under the null-hypothesis of Gaussian white-noise, it has the nature of the ratio of two independent statistics and it behaves as a Snedecor F statistic with 2 and degrees of freedom. It is reminiscent of a ratio of two quadratic forms, similarly to the use of nested models (here a pure cosine model versus a constant term). Since the global power is proportional to the variance of the white-noise process, it is mandatory to normalise the power by a function of this variance. This is precisely the purpose of the statistic.
@@
The periodogram is explored and the frequencies corresponding to the various peaks that are above a certain threshold are further retained. This threshold is defined as the limit above which the highest peak can be considered as significant at a certain probability level. This is further described in the next Section. We keep in a list up to one hundred candidate frequencies. This is more than the amount of independent ones in the classical Fourier. If too many peaks are listed, only the 100 highest ones are kept. For each of these selected candidate frequencies, a fit is made of an eccentric Keplerian orbital model. At this stage, there is a possibility to add one additional candidate frequency coming for example from CU7 (e.g. on the basis of a photometric curve). In DR3 in particular, this possibility is only used for the subset of solutions of the type EclipsingSpectro combined with SB1 solutions (see section 7.6 by Siopis & Sadowski in Pourbaix et al. 2022; Siopis et al. 2024).

### v2

@@ 3 The underlying orbital models
Four kinds of orbital solutions are considered here: they are labelled SB1, SB1C, TrendSB1, and StochasticSB1. They are described here below. The single-line RVs treated here and consequently the deduced motion is intended to refer to the sole object that is visible in the spectra. The model will be fitted through a classical least-square procedure and the selected objective function is the -function except when otherwise stated. The uncertainties on the parameters are calculated in the classical way on the basis of the diagonal elements of the inverse of the curvature matrix (i.e. the variance-covariance matrix; see Bevington & Robinson 2003). Despite the fact that we use the Greek symbol for all (i.e. concerning input data as well as output parameters) uncertainties, it should be pointed out that all of them are of course estimators. They are always expressed as 1 values.
@@ 3.1 Orbital models of type SB1
Here, is the time of passage at periastron, while is the argument of periastron measured in the sense of the orbit with its origin at the ascending node. In this model, the eccentricity enters as a strongly non-linear parameter. This is also the case for . The parameters to be determined are , , , , , (respectively, the period, the centre-of-mass velocity, the semi-amplitude, the eccentricity, the argument of periastron and the time of passage at periastron). The first step of the computation is devoted to the determination of the period. Some significance tests are made on the very existence of the period. If the period is not significant, the model is fitting noise and is considered as being invalid. This is a statistical decision. Owing to the insufficient span of time covered for the DR3, no apsidal motion is considered; is considered constant in time.
@@ 3.3 Models of type TrendSB1
It can happen that the RV evolution as a function of time just exhibits a trend. This could be a transient behaviour (e.g. outburst) just observed by chance or correspond to a small piece of slowly accelerated motion (e.g. of a stellar atmosphere). It could also correspond to an orbital motion with a period well longer than the span of observing time covered by the satellite. This span of time is 34 months for the DR3. However, it could be slightly shorter for some objects because it combines with the gaps in the scanning law. The present modelisation is restricted to polynomials of one to four degrees and presenting only one extremum on the interval. The degree of the polynomial is chosen on the basis of an F-test for nested models. In practice, it turned out that only the one- and two-degree polynomials were necessary. No higher degree fit was required. The formalism used for the fit is the same as for Hipparcos (ESA 1997). The TrendSB1 solution is compared to the orbital solutions using a Fisher Snedecor F-test on the comparison of the respective s. It should be made clear that this comparison could be an approximation.
@@ 3.5 The downstream combination of solutions
The outputs of the NSS spectroscopic pipeline are delivered to the combiner that operates downstream. The task of the combiner is to inspect the output of the astrometric, photometric, and spectroscopic NSS channels in order to detect possible solutions corresponding to the same object and that could be combined. This will be described in Sect. 7.2.
@@ 4 Determination of the periodicity
The first step in the analysis is the search for a periodic behaviour of the RVs. The analysis is based on a periodogram (actually a frequencygram), which has the nature of the Fourier power spectrum. Since the time sampling of the Gaia satellite is particularly complex, the classical Fourier formalism is not applicable. To aptly understand the limitation of the method, it is necessary to have some knowledge of the impact of the sampling and of a few characteristic timescales.
@@ 4.1 Effect of the Gaia sampling
To explain the effect of the Gaia sampling, we have to start from a single, idealistic, case and progressively add certain degrees of complexity. We start by considering the spinning of the satellite on itself that makes that a star is measured every 0.25 d (this is our first basic idealistic case since this simple view is never occurring in practice). The existence of the sample periodicity concerns both fields of view individually. Therefore, 0.25 d is the typical, smallest regular and periodic timescale in the sampling. The 0.25 d step should generate in the time series some aliasing with a typical 2 d-1 Nyquist frequency. This means that a progenitor frequency will be duplicated at , , , and so on (in d-1 units). The duplication is also extended to the negative frequencies. This duplication represents a pure aliasing, since the sampling is in this hypothetical case regular. Therefore, the aliasing is pure and the Nyquist frequency exists. In such a case, intermediate and large periods could have aliases towards periods of 0.25 to 0.5 d (and vice versa).
@@
However, the existence of two fields of view separated by 1065 on the sky renders the situation more complex by adding a regularity that is not properly speaking periodic. Thus, the aliasing is no more pure, we will call it pseudo-aliasing. The semi-amplitude spectral window is a tool that facilitates the analysis of such phenomenon. It has been introduced by Deeming (1975, 1976). In the case of an equidistant sampling, the spectral window is the well-known Dirac comb (or Shah function; see Bloomfield 1976). The general spectral window represents an extension of this Dirac comb to the general case of odd sampling. The property that the generalised Fourier Spectrum of the sampled data is the result of the convolution of the underlying Fourier Spectrum of the signal by the spectral window remains true for almost any kind of sampling (Deeming 1975, 1976). If the satellite was limited to a rotation on itself, the semi-amplitude spectral window corresponding to the combined two fields of view would resemble the plot in the upper panel of Fig. 1. Contrary to the classical case of pure aliasing, the aliasing is here not pure since all the peaks have different heights. Besides the main peak at , the next first peak is at d-1, but it is mildly high. The peak having a height comparable to the peak is the one at d-1 which could represent some almost pure aliasing. This figure represents an idealistic case. First, it is assumed that the data points in each field of view are perfectly regularly distributed. This is not the case because several phenomena induce a small perturbation of the time of the observations. The variations of the Basic Angle as well as micro-clanks and meteoritic-hits are perturbing the regularity (Gaia Collaboration et al. 2016b). We do not further consider these effects in our explanation. The satellite is not only spinning on itself, but also follows a more complicated motion. First, the rotation axis is precessing along an axis directed towards a fictitious nominal position of the Sun (compensating for the orbit of the satellite around the Sun-Earth Lagrangian point L2). In addition, the satellite located around the Lagrangian point L2 is orbiting around the Sun with the Earth annual motion (change in ecliptic longitude of the Sun, Gaia Collaboration et al. 2016b). The major consequence is that, for a given object, the majority of the turns of the spinning satellite will not correspond to a transit of the object and thus no observation will be acquired for this object. Most of the observations are not done and medium and large gaps occur during which the field of view is not directed towards the star. This effect is dependent on the position of the object on the sky (Gaia Collaboration et al. 2016b). Therefore, the sky sampling is markedly non-uniform (see Sect. 5.2 of Gaia Collaboration et al. 2016b).
@@
Practically, this generates in the spectral window a power leakage from the main peak at d-1 towards other frequencies. This effect is illustrated in the other panels of Fig. 1. We produce typical time series of Gaia RVS with the GOST tool222https://gaia.esac.esa.int/gost/. Fig. 1 second panel represents a typical spectral window of a time series containing 26 observations, the third panel 19 and the bottom one 14 observations (against the observations for the first panel). This small number of observations induces an aliasing but the aliasing typical of the Gaia sampling is not leaking power towards a very few other frequencies but towards numerous frequencies. Besides the main peak, there is a leakage towards a forest of frequencies. The contrast between the parasitic peaks and the main one is decreasing with the number of data points (following a for the semi-amplitude). Typically, pseudo-aliasing can occur at any . If existing, the value of is most probably located below let say 3 d-1. The structure due to the motion considered in the top panel of Fig. 1 almost totally disappeared. In the power spectrum of the signal, the combination of noise and of aliasing is going to generate parasitic peaks at frequencies that mainly depend on the position of the object on the sky. Consequently, the time sampling is complex, and the aliasing is dependent on the star being observed.
@@
Looking at the third panel of Fig. 1, we notice a marked aliasing at and at . Therefore, in a few cases, noise could conspire to shift the largest peak position from the progenitor peak by this amount, thus producing pseudo-aliases of the progenitor that could dominate. However, in the present case, predominant aliasing seems to be restricted to small , i.e. to small shifts of the progenitor peak. Other values of will be associated (or not) to other regions of the sky. Large shift aliasing is clearly possible but rather for values . This should have no impact in the framework of spectroscopic binaries.
@@
Here above, we have shown three examples illustrating the strong leakage characterising the Gaia sampling (for RVS). An alternative view would be to compute some kind of typical spectral window. We obtain such a result using simulations. We generated 65 536 positions randomly and uniformly distributed over all the sky. For each position, we extracted from the GOST software the series of observational epochs typical of this particular position and corresponding to the nominal scanning law (NSL) for RVS. On purpose, we discarded the first month of scanning which was characterised by the ecliptic pole scanning law (EPSL). For the latter, the precession of the satellite was stopped and the scanning law clearly becomes more relevant to the scanning due to the spinning of the satellite, except for a slow rotation due to the orbital motion around the Sun. We thus neglected the EPSL to instead concentrate on the effect of the NSL. For each simulation, we derived the corresponding amplitude spectral window (for the RVS instrument). Figure 2 illustrates (in red) the median of the amplitude spectral windows. Despite the fact that we restricted our simulations to the NSL, the aliasing due to the spinning of the satellite is clearly visible. However, even the peaks at d-1 (and at d-1) are rather low, compared to the peak at d-1, at most 30% (in amplitude). We also plot in Fig. 2 (in black) 32 randomly selected amplitude spectral windows. Clearly, the dispersion of the various realisations of the amplitude spectral window is rather large. Also, near the possible alias frequencies, the amplitude spectral window could present values nearing zero. This confirms the fact that in most cases, the spinning pseudo-aliasing will have little, if any, impact on the global aliasing. Although a few positions on the sky could still exhibit the effect of the spinning aliasing. Of course, if, over the 34 months, the EPSL scanning is dominating the definition of the times in the analysed time series, the effect will increase exhibiting a stronger spinning aliasing. We must insist here on the fact that no signal is analysed in the present section; the dispersion of the amplitude spectral windows is originating from the randomness of the selection of the sky positions and from the variety of nature of the aliasing as a function of the latter. For any time series, the observing time could be used to pre-calculate the particular spectral window independently of the signal.
@@ 4.2 Search for the periodicity
When a periodicity is present in a time series, the periodogram exhibits a peak at the corresponding frequency (the progenitor), but other peaks are also present at the location of the pseudo-aliases of the progenitor. Contrary to the case of even sampling, the actual sampling generates pseudo-aliases that are in principle lower than the progenitor peak. However, the combination of the aliasing with background noise or other phenomena (blending) induces the fact that the choice of the true progenitor among the aliased peaks is never fully secure.
@@
The total span of time of the DR3 time series is at most of 34 months and thus the largest periodicity accessible from the point of view of the Fourier corresponds roughly to a frequency of 1/1000 d-1. This value could actually be larger taking into account the gaps generated by the actual sampling law. The constraints should thus be expressed as a function of the actual length of the time series and thus as a function of . It has been decided that it would be adequate to limit the range of investigation for periodicities to since beyond the possible orbit is certainly not well determined. However, from the practical point of view, we decided to limit the range for period search to although no orbit are computed for periods located in this extension (see Sect. 6.3 for more details). The adopted frequency step is actually where is the actual total time-span of the individual time series. This represents an oversampling of a factor of 100, which gives direct access to the amplitude of the signal as the height of the peak. Based on the discussion detailed in the previous section, we decided that the adopted largest frequency to be explored is 4 d-1. This choice is rather atypical since it was chosen to correspond to the spin period.
@@
If the star indeed varies in a periodic way, the largest peak in the periodogram is an estimate of the possible period, but does not stand as definitive proof of its existence. We should consider that a periodicity is proven if a minimum of three identical cycles are observed. Although the long periods are indicative, they are associated to large errors due to the small number of cycles included in the computation.
@@
The method used here is that of Heck-Manfroid-Mersch (hereafter HMM: Heck et al. 1985; Gosset et al. 2001), which advantageously replaces the classical Fourier method (see e.g. Bloomfield 1976) in the case of odd sampling. In principle, the highest peak is assumed to be the progenitor one and to correspond to the orbital period. This last point is particularly true for the case of RV orbital motion, since most of the power is gathered in the fundamental frequency; the power of the peaks corresponding to high harmonics is gradually decreasing in the case of the Keplerian motion. This is due to the fact that RV curves always present one simple cycle during the orbital period. On the contrary, this is usually not the case for light curves of eclipsing binaries and ellipsoidal variables where most of the power is gathered in twice the orbital frequency, corresponding to the average depth of the minima, whereas the power at the orbital frequency is dominated by the difference between the primary and secondary minima and thus is much weaker. In their paper, HMM introduced the statistic which is a power related statistic (see Eq. A.11 in Gosset et al. 2001). At a predetermined frequency and under the null-hypothesis of Gaussian white-noise, it has the nature of the ratio of two independent statistics and it behaves as a Snedecor F statistic with 2 and degrees of freedom. It is reminiscent of a ratio of two quadratic forms, similarly to the use of nested models (here a pure cosine model versus a constant term). Since the global power is proportional to the variance of the white-noise process, it is mandatory to normalise the power by a function of this variance. This is precisely the purpose of the statistic.
@@
The periodogram is explored and the frequencies corresponding to the various peaks that are above a certain threshold are further retained. This threshold is defined as the limit above which the highest peak can be considered as significant at a certain probability level. This is further described in the next section. We keep in a list up to 100 candidate frequencies. This is more than the amount of independent ones in the classical Fourier. If too many peaks are listed, only the 100 highest ones are kept. For each of these selected candidate frequencies, a fit is made of an eccentric Keplerian orbital model. At this stage, there is a possibility to add one additional candidate frequency coming for example from CU7 (e.g. on the basis of a photometric curve). In DR3 in particular, this possibility is only used for the subset of solutions of the type EclipsingSpectro combined with SB1 solutions (see Sect. 7.6 by Siopis & Sadowski in Pourbaix et al. 2022; Siopis et al. 2024).